a canoeist Paddled For 4 Hours With A 6 Km/h Current To Reach A Campsite. The Return Trip Against The Current Took 10 Hours. Find The Speed Of The Boat In Still Water?
Answer by mananth(16946)

a canoeist Paddled For 4 Hours With A 6 Km/h Current To Reach A Campsite. The Return Trip Against The Current Took 10 Hours. Find The Speed Of The Boat In Still Water?
...
speed of boat in still water
..
with current speed = x+6: time = 4 hours
against , speed = x-6: time = 10 hours
...
distance = speed * time
4(x+6)=10(x-6)
4x+24=10x-60
6x=84
x=14 km/h speed of boat
...
m.ananth@hotmail.ca